# From Prejudice to Parity: A New Approach to Debiasing Large Language Model Word Embeddings

### v2

@@ Abstract
Embeddings play a pivotal role in the efficacy of Large Language Models. They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform remarkably on a plethora of complex tasks that require a fundamental understanding of human language. Given that these embeddings themselves often reflect or exhibit bias, it stands to reason that these models may also inadvertently learn this bias. In this work, we build on the seminal previous work and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’. We exhaustively evaluate this algorithm across a variety of SOTA datasets, accuracy metrics, and challenging NLP tasks. We find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
@@ 1 Introduction
Word embeddings are a foundational element in the architecture of large language models (LLMs). They act as the basis for these models to understand and subsequently, generate human-like language. However, it has been shown that these models rely on word embeddings, which themselves may reflect or exhibit bias Papakyriakopoulos et al. (2020); Dev et al. (2020). Given the exponential increase in the use of large language models on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice. Existing work in this field has looked most commonly into gender bias Kotek et al. (2023); Bordia and Bowman (2019); de Vassimon Manela et al. (2021), racial bias Mozafari et al. (2020); Omiye et al. (2023); Tang et al. , and religious bias Baligudam (2022); Kirk et al. (2021). In this work, we build on the seminal work of Gonen and Goldberg (2019), which brought attention to the inherent biases present in traditional GloVe embeddings Pennington et al. (2014). This study prompted the NLP community to reevaluate the fundamental choices underlying our word representation models. Specifically, we present DeepSoftBias: an algorithm that furthers the application of their methodology, by diverging from the conventional GloVe embeddings and delving into the word embeddings produced by the best-performing models on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2022) leaderboard. By employing these advanced embeddings on the same set of words as used in GloVe embeddings, we seek to investigate whether these state-of-the-art (SoTA) models inherently exhibit reduced bias. Our primary objective is twofold: first, to de-bias the embeddings from these selected models, and second, to rigorously assess the effectiveness of the bias removal process. Our proposed approach, DeepSoftDebias, is an innovative methodology to de-bias LLM word embeddings which involves integrating a neural network into the soft debiasing approach developed by Bolukbasi et al. (2016). This novel amalgamation is driven by the aspiration to enhance the debiasing process and contribute to the ongoing discourse on creating fair and ethically sound language models. To this end, our work answers the following research questions:
@@
Our Contributions
[leftmargin=1mm] We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion. We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines
@@ 2 Data
2.1 L2-Reddit Corpus
We utilize the L2-Reddit [^1] https://github.com/ellarabi/reddit-l2 Rabinovich et al. (2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences from the United States. This dataset serves as the foundation for training the word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse. The corpus provides a rich linguistic resource for analyzing and understanding the initial biased word embeddings. In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al. (2013) word embeddings. Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the Language Model Models (LLMs).
2.2 StereoSet
StereoSet Nadeem et al. (2020) stands out as a critical dataset for measuring stereotype bias in language models, containing around 17,000 sentences across demographic dimensions like gender, race, religion, and profession. It introduces the Context Association Tests (CAT) for evaluating model preferences and biases, providing a structured approach to assess and quantify biases in popular models like BERT Devlin et al. (2019), GPT-2 Radford et al. (2019), RoBERTa Liu et al. (2019), and XLNet Yang et al. (2020). The dataset’s creation involved a rigorous crowdsourcing process, ensuring a wide coverage of stereotypes and anti-stereotypes to challenge and evaluate the fairness of NLP systems. In our words we use the stereoset dataset to benchmark our debiasing method against all the other debiasing methods. To accomplish this, we compute the average sentence vector for all sentences and their respective contexts. We then compare these vectors against each other to derive the stereotype score.
2.3 CrowS-Pairs
CrowS-Pairs Nangia et al. (2020), designed to assess social biases in masked language models (MLMs), comprises 1,508 examples covering nine bias types, including race, religion, and age. It contrasts sentences related to historically disadvantaged and advantaged groups in the U.S., with annotations from crowd workers highlighting the degree of stereotyping. Despite its noise and reliability issues, CrowS-Pairs remains a valuable resource for evaluating the bias in major language models, providing a nuanced metric that quantifies preference for stereotypical versus non-stereotypical expressions. In our study, we obtain debiased word embeddings for sentences by computing the average sentence vector for both less and more stereotypical or anti-stereotypical directions. We then compare these embeddings against each other to calculate the Crows Metric score.
@@ 3 Methodology
In this research, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations. These biases span across gender, racial, and religious lines and are encoded inadvertently within language models.
@@ 3.1 Bias Identification and Data Structure
To quantitatively assess bias in word embeddings, we measure the projection of word vectors onto a gender-specific axis, defined by the vectorial difference between the terms ‘he’ and ‘she.’ The magnitude of this projection serves as an indicator of bias. We use a structured vocabulary with its associated vector representations from the Word2Vec model to facilitate the identification of biases. For a comprehensive evaluation, we utilize additional data files that include definitive sets of gender-associated word pairs, analogy templates that list occupational roles often linked with specific genders, and a set of neutral terms used as evaluation targets. These resources are crucial for the systematic identification and rectification of biases in word embeddings.
@@ 3.2 Soft Debiasing: The Baseline Approach
The initial method as seen in the paper Manzini et al. (2019) leverages a method called soft debiasing. We recap its algorithm in Algorithm 1. Soft debiasing involves learning a projection of the embedding matrix that preserves the inner product between biased and debiased embeddings while minimizing the projection onto the bias subspace of embeddings mentioned in 3.1 that should be neutral. Given embeddings $W$ and $N$ which are embeddings for the whole vocabulary and the subset of bias-neutral words respectively, and the bias subspace B obtained in Section 2.1, soft debiasing seeks a linear transformation A that minimizes the following objective defined in Eq. (1) as follows:
@@
Minimizing the first term preserves the inner product after the linear transformation A, and minimizing the second term minimizes the projection onto the bias subspace B of embeddings. R is a tunable parameter that balances the two objectives.
@@ 3.3 DeepSoftDebias: Our Proposed Approach
In the original approach introduced by Bolukbasi et al. (2016), a transformation matrix is utilized and optimized by an optimizer to enable a direct mapping between input and output embeddings. To enhance performance, we propose DeepSoftDebias. In this approach, we replace the transformation matrix with a neural network, leveraging its capability to represent a sequence of transformation matrices. This adaptation enables the algorithm to handle more complex functions mapping between input and output embeddings. We use the same loss functions as mentioned in the section 3.2. Furthermore, we transition from stochastic gradient descent (SGD Robbins and Monro (1951)) to the Adam Kingma and Ba (2017) optimizer, resulting in enhanced efficiency, speed, and optimization quality. We describe our full algorithm in Algorithm 2. While these modifications were implemented, the fundamental aspects of the method remain unaltered, ensuring minimal alterations in embeddings and preserving orthogonality with the bias space. Unlike the baseline, which relies on singular value decomposition (SVD) and incurred information loss, DeepSoftDebias preserves the full information of the original matrix. Moreover, unlike the baseline, DeepSoftDebias can handle large embedding dimensions of more than 4.5k. We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency. Fig. 2 presents a visualization of our approach to downstream testing.
@@
Figure 2: A step-by-step visualization of our downstream testing process to effectively evaluate DeepSoftDebias
@@ 4 Effects of LLM Size and Dependency of Network Size
The debiasing performance of word embeddings depends on the size of the embeddings and the depth of the debiasing neural network, rather than the number of parameters of the language model. We observe in 11 Smaller models, such as bge-small Xiao et al. (2023) and DeBERTa-v3-base He et al. (2023) or DeBERTa-v3-large, can be debiased effectively by a single-layer neural network. Larger models, such as Llama-2 Touvron et al. (2023), Alpaca Taori et al. (2023) and Yi-6b 01.ai (2024) need a more complex debiasing neural network. For embedding dimensions around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results. Additionally, the debiasing neural network and the optimization algorithm need to be hyperparameter-tuned, such as adjusting the learning rate, to get optimal results. The hyperparameters may vary depending on the model size, the embedding dimension, and the debiasing task. Therefore, a systematic search for the best hyperparameters is necessary to ensure the effectiveness of the debiasing process.
@@ 5 Results
The results obtained are described in the following subsections.
5.1 Mean Average Cosine Similarity (MAC)
MAC Manzini et al. (2019) is a metric used to quantify semantic associations between word classes and attributes. MAC takes word embeddings, targets (representing classes), and attributes as inputs. By computing the mean cosine distance between target words and attribute sets, MAC offers a concise measure of semantic proximity. This metric provides valuable insights into the contextual semantics encoded within word embeddings. Table 1 shows that the word embeddings debiased in the direction of race and gender have comparable increases in their average MAC of  0.64, whereas word embeddings debiased in the direction of religion have an increase in MAC of around 0.61. We see that our debiasing procedure categorically moves MAC scores closer to 1.0. This indicates an increase in cosine distance. Further, the associated P-values indicate these changes are statistically significant. This demonstrates that our approach for multiclass debiasing decreases bias in the word embeddings. We provide visual representations of the efficiency of DeepSoftDebias at removing gender bias, racial bias, and religion bias in Appendix A.

### v4

@@ Abstract
Embeddings play a pivotal role in the efficacy of large language models. They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform complex tasks that require a fundamental understanding of human language. Given that these embeddings themselves often reflect or exhibit bias, it stands to reason that these models may also inadvertently learn this bias. In this work, we build on the aforementioned seminal work of Bolukbasi et al. (2016) and Gonen and Goldberg (2019) and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’. We exhaustively evaluate this algorithm across a variety of state-of-the-art datasets, accuracy metrics, and challenging NLP tasks. We find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
@@ 1 Introduction
Word embeddings are a foundational element in the architecture of Large Language Models (LLMs). They act as the basis for these models to understand and subsequently, generate human-like language. However, it has been shown that these word embeddings themselves may reflect or exhibit bias Dev et al. (2020); May et al. (2019); Caliskan et al. (2017). Given the exponential increase in the use of LLMs on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice. Existing work in this field has looked most commonly into gender bias Kotek et al. (2023); Bordia and Bowman (2019); de Vassimon Manela et al. (2021), racial bias Mozafari et al. (2020); Omiye et al. (2023); Tang et al. , and religious bias Baligudam (2022); Kirk et al. (2021). In this work, we build on the seminal work of Gonen and Goldberg (2019), which brought attention to the inherent biases present in traditional GloVe embeddings Pennington et al. (2014). This study prompted the NLP community to reevaluate the fundamental choices underlying our word representation models. Specifically, we present DeepSoftBias: an algorithm that furthers the application of their methodology, by diverging from the conventional GloVe embeddings and delving into the word embeddings produced by the best-performing models on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2022) leaderboard. By employing these advanced embeddings on the same set of words as used in GloVe embeddings, we seek to investigate whether these state-of-the-art (SoTA) models inherently exhibit reduced bias.
Our primary objective is twofold: first, to de-bias the embeddings from these selected models, and second, to rigorously assess the effectiveness of the bias removal process. Our proposed approach, DeepSoftDebias, is an innovative methodology to de-bias LLM word embeddings which involves integrating a neural network into the soft debiasing approach developed by Bolukbasi et al. (2016). This novel amalgamation is driven by the aspiration to enhance the debiasing process and contribute to the ongoing discourse on creating fair and ethically sound language models. To this end, our work answers the following research questions:
@@
Our Contributions We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion. We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines.
@@
2 Related Work
INLP
Iterative Null-space Projection (INLP) Ravfogel et al. (2020) is a post-hoc debiasing method that operates at the representation level. The INLP methodology debiases representations by iteratively projecting them into a linear classifier’s null space. This technique is particularly effective for handling intersectional groups, which are defined by combinations of sensitive attributes². INLP seeks to learn a hidden representation that is independent of the protected attributes. This approach is beneficial in scenarios where an attempt to make a model fairer towards some group results in increased unfairness towards another group. Therefore, INLP emerges as a robust and effective strategy for mitigating bias in language models, promoting fairness across multiple protected attributes.
Self-Debias
Self-Debiasing Schick et al. (2021) is a novel approach to mitigating bias in language models. The methodology, first coined by Schick et al. (2021), is based on the concept of self-diagnosis. In this approach, pretrained language models recognize their undesirable biases and the toxicity of the content they produce. Based on this self-diagnosis, a decoding algorithm is proposed that reduces the probability of a language model producing problematic text. This approach, referred to as self-debiasing, does not rely on manually curated word lists, nor does it require any training data or changes to the model’s parameters. While it does not completely eliminate the issue of language models generating biased text, it is an important step in this direction. The self-debiasing approach demonstrates the potential of language models to self-regulate and reduce their inherent biases.
Sentence Debias
SentenceDebias Liang et al. (2020) is a debiasing methodology that operates at the sentence level. It is a projection-based method that identifies a linear subspace associated with a specific bias. The sentence representations are projected onto this bias subspace, and the projection is subtracted from the original representations. This process effectively debiases the sentence representations. SDB is particularly useful for mitigating biases related to gender, race, and religion. It offers a comprehensive comparison between models that adjust weights for debiasing and those employing test-time surgical interventions. The SDB method signifies a significant advancement in debiasing strategies, promoting a more equitable representation in language models.
Counterfactual Data Augumentation
Counterfactual Data Augmentation (CDA) Yadav et al. (2023) is a data-based debiasing strategy often used to mitigate gender bias. The CDA methodology involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset. This technique is part of a broader set of debiasing techniques that also includes Dropout, Self-Debias, SentenceDebias, and Iterative Nullspace Projection. CDA has been applied to various language models, including BERT, with the goal of diminishing stereotypical biases while maintaining the model’s performance on downstream tasks. However, it’s important to note that while CDA has the potential to improve the fairness of NLP models, it may not be effective in eliminating all biases and may even introduce new biases or errors in the model³.
FineDeb
FineDeb Saravanan et al. (2023) is a two-phase debiasing framework for language models. In the first phase, FineDeb debiases the model by modifying the embeddings learned by the language model. This process involves contextual debiasing of these embeddings. In the second phase, the debiased model is fine-tuned on the language modeling objective. This methodology is effective for demographics with multiple classes. The FineDeb approach demonstrates its effectiveness through extensive experiments and comparisons with state-of-the-art techniques. It offers stronger debiasing in comparison to other methods, which often result in models as biased as the original language model. Thus, FineDeb emerges as a robust and effective framework for mitigating bias in language models.
@@ 3 Data
L2-Reddit Corpus
We utilize the L2-Reddit [^1] https://github.com/ellarabi/reddit-l2 Rabinovich et al. (2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences. This dataset serves as the foundation for training word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse. In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al. (2013) word embeddings. Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the LLMs. We utilize Word2Vec on the Reddit-L2 corpus to obtain the vocabulary. This vocabulary comprises the words for which we aim to extract embeddings from the LLMs. The primary objective of this approach is to ensure a consistent set of words across all our LLMs. This consistency allows each of our LLMs to be tested on the same set of words.
StereoSet
StereoSet Nadeem et al. (2020) stands out as a critical dataset for measuring stereotype bias in language models, containing around 17,000 sentences across demographic dimensions like gender, race, religion, and profession. It introduces the Context Association Tests (CAT) for evaluating model preferences and biases, providing a structured approach to assess and quantify biases in popular models like BERT Devlin et al. (2019), GPT-2 Radford et al. (2019), RoBERTa Liu et al. (2019), and XLNet Yang et al. (2020). In our work, we use the Stereoset dataset to benchmark our debiasing method.
CrowS-Pairs
CrowS-Pairs Nangia et al. (2020), designed to assess social biases in masked language models (MLMs), comprises 1,508 examples covering nine bias types, including race, religion, and age. It contrasts sentences related to historically disadvantaged and advantaged groups in the U.S., with annotations from crowd workers highlighting the degree of stereotyping. In our study, we obtain debiased word embeddings for sentences by computing the average sentence vector for both less and more stereotypical or anti-stereotypical directions. We then compare these embeddings against each other to calculate the Crows Metric score.
@@ 4 Methodology
In this section, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations. These biases span across gender, racial, and religious lines and are encoded inadvertently within language models.
@@ 4.1 Bias Identification and Data Structure
To quantitatively assess bias in word embeddings, we measure the projection of word vectors onto a gender-specific axis, defined by the vector difference between the terms ‘he’ and ‘she.’ The magnitude of this projection serves as an indicator of bias. We use a structured vocabulary with its associated vector representations from the Word2Vec model to facilitate the identification of biases. For a comprehensive evaluation, we utilize additional data files that include definitive sets of gender-associated word pairs, analogy templates that list occupational roles often linked with specific genders, and a set of neutral terms used as evaluation targets. These resources are crucial for the systematic identification and rectification of biases in word embeddings. The words used for the BiasSpace are present in AppendixA.
@@ 4.2 Soft Debiasing: The Baseline Approach
The initial method as seen in Manzini et al. (2019) leverages a method called soft debiasing. We recap its algorithm in Algorithm 1. Soft debiasing involves learning a projection of the embedding matrix that preserves the inner product between biased and debiased embeddings while minimizing the projection onto the bias subspace of embeddings mentioned in 4.1. Given embeddings $W$ and $N$ which are embeddings for the whole vocabulary and the subset of bias-neutral words respectively, and the bias subspace B obtained in Section 3, soft debiasing seeks a linear transformation A that minimizes the following objective defined in Eq. (1) as follows:
@@
Minimizing the first term preserves the inner product after the linear transformation A, and minimizing the second term minimizes the projection onto the bias subspace B of embeddings. $\lambda$ is a tunable parameter that balances the two objectives. W here refers to the matrix of word embeddings and N refers to the matrix of the embeddings of the neutral space i.e. words that aren’t influenced by any bias.
@@ 4.3 DeepSoftDebias: Our Proposed Approach
In the original approach introduced by Bolukbasi et al. (2016), a transformation matrix is utilized and optimized by an optimizer to enable a direct mapping between input and output embeddings. To enhance performance, we propose DeepSoftDebias. In this approach, we replace the transformation matrix with a neural network, leveraging its capability to represent a sequence of transformation matrices. This adaptation enables the algorithm to handle more complex functions mapping between input and output embeddings. We use the same loss functions as mentioned in the section 4.2. Furthermore, we transition from stochastic gradient descent (SGD Robbins and Monro (1951)) to the Adam Kingma and Ba (2017) optimizer, resulting in enhanced efficiency, speed, and optimization quality. We describe our full algorithm in Algorithm 2. While these modifications were implemented, the fundamental aspects of the method remain unaltered, ensuring minimal alterations in embeddings and preserving orthogonality with the bias space.
Unlike the baseline, which relies on singular value decomposition (SVD) and incurred information loss, DeepSoftDebias preserves the full information of the original matrix. Moreover, unlike the baseline, DeepSoftDebias can handle large embedding dimensions of more than 4.5k. We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency. The reason for the need of a fixed Biasspace is that we adopt the methodology proposed by Bolukbasi et al. for the derivation of the bias subspace.
The process of creating the BiasSpace commences with the identification of word vectors representing opposing concepts, such as ‘he’ versus ‘she’, or ‘man’ versus ‘woman’. For each pair, we compute the mean vector, which encapsulates the shared semantic space. Subsequently, we subtract this mean vector from the original word vectors, yielding vectors that exclusively represent the bias components. These bias vectors are then concatenated to form a matrix, referred to as the bias subspace. This bias subspace plays a pivotal role in the training of our neural network. Specifically, we ensure that the output of the word embeddings, upon being processed through the neural network, is orthogonal to the bias subspace Fig. 2 presents a visualization of our approach to downstream testing.
@@
Figure 2: A step-by-step visualization of our downstream testing process to effectively evaluate DeepSoftDebias.
@@ 5 Effects of LLM Size and Dependency of Network Size
The debiasing performance of word embeddings depends on the size of the embeddings and the depth of the debiasing neural network, rather than the number of parameters of the language model. We observe in 11 Smaller models, such as bge-small Xiao et al. (2023) and DeBERTa-v3-base He et al. (2023) or DeBERTa-v3-large, can be debiased effectively by a single-layer neural network. Larger models, such as Llama-2 Touvron et al. (2023), Alpaca Taori et al. (2023) and Yi-6b 01.ai (2024) need a more complex debiasing neural network. For embeddings with embedding length of around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results. In addressing the second research question, we delve into the intricacies of neural network complexity necessary for debiasing embeddings of varying sizes. While our discussion highlights the effectiveness of larger neural networks in mitigating bias within Language Model (LM) embeddings with substantial dimensions, it is imperative to substantiate this observation. We would like to point out that we draw inspiration from the conceptual framework of DeepSoftDebias. Building upon the foundational work by Bolukbasi et al., which employed a transformation matrix for word embedding debiasing, our approach replaces this matrix with a neural network. This neural network can be conceptualized as a series of interconnected matrices. Specifically, when de-biasing larger LMs with embedding dimensions exceeding 4096, we augment the neural network by increasing the number of layers and adjusting layer sizes. This augmentation enables us to model the intricate dependencies inherent in debiasing processes for larger embedding dimensions. Consequently, deeper neural networks emerge as more efficacious tools for addressing bias in such expansive models. Additionally, the debiasing neural network and the optimization algorithm need to be hyperparameter-tuned, such as adjusting the learning rate, to get optimal results. The hyperparameters may vary depending on the model size, the embedding dimension, and the debiasing task. Therefore, a systematic search for the best hyperparameters is necessary to ensure the effectiveness of the debiasing process.
@@ 6 Results
In this section, we provide an extensive analysis of our proposed methodology, complete with a comprehensive evaluation against multiple metrics, tasks, and datasets. We provide the results of additional downstream testing and ablation experiments in Appendix D and Appendix F, respectively. We also provide our hypothesis of why there is a variation in bias across LLMs in Appendix E.
@@
Gemma Gemma-2b 0.058 0.971 47.107 48.470 57.69 36.27
Gemma Gemma-7b 0.553 0.976 49.173 51.53 63.46 35.29
GritLM GritLM-7B 0.379 0.999 51.239 48.470 57.05 37.25
mxbai mxbai-embed-large-v1 0.467 0.994 51.652 55.34 60.9 49.02
@@
6.1 Mean Average Cosine Similarity
Mean Average Cosine Similarity (MAC) Manzini et al. (2019) is a metric used to quantify semantic associations between word classes and attributes. MAC takes word embeddings, targets (representing classes), and attributes as inputs. By computing the mean cosine distance between target words and attribute sets, MAC offers a concise measure of semantic proximity. This metric provides valuable insights into the contextual semantics encoded within word embeddings. Table 1 shows that the word embeddings debiased in the direction of race and gender have comparable increases in their average MAC of 0.64, whereas word embeddings debiased in the direction of religion have an increase in MAC of 0.61. We see that our debiasing procedure categorically moves MAC scores closer to 1.0. This indicates an increase in cosine distance. Further, the associated P-values indicate these changes are statistically significant. This demonstrates that our approach for multiclass debiasing decreases bias in the word embeddings. We provide visual representations of the efficiency of DeepSoftDebias at removing gender bias, racial bias, and religion bias in Appendix B.
In our research, we have chosen to utilize Mean Average Cosine Similarity (MAC) as our primary metric for assessing bias in word embeddings. This decision is informed by the work of Manzini et al. (2019), who posit that MAC can be viewed as an extension of the Word Embedding Association Test (WEAT), specifically adapted for a multiclass setting. The MAC and WEAT serve distinct, yet complementary purposes. While WEAT is designed to focus on specific associations between word vectors and predefined concepts (such as gender or race), MAC provides a broader perspective by measuring overall similarity patterns across different groups. This makes MAC less sensitive to specific word choices, thereby revealing biases that might be overlooked by WEAT. In essence, both metrics contribute to a comprehensive understanding of bias in word embeddings. However, the use of MAC is particularly beneficial in our research as it complements the findings of WEAT, providing a more holistic view of bias in the data. This approach allows us to capture a wider range of biases, thereby enhancing the robustness of our analysis.